Oatmeal often comes in a container shaped like a cylinder. Suppose the volume V of an oatmeal container is 4,618.1 cubic centimeters. Solve the equation
V = pi r^2 h for h. Then find the height h of the container if the radius is 7 centimeters. Round to the nearest tenth.

Answer by ewatrrr(24785)


Hi
V(cylinder) = pi*r^2 h |Note the use of ^ (uppercase 6)
r = 7cm V = 4618.1 cm^3
V(cylinder) = pi*r^2 h
V/pi*r^2 = h
4618.1/(3.14*49) = h
30.0cm = h (rounded to the nearest tenth)